# Evaluating the Defense Potential of Machine Unlearning against Membership Inference Attacks

### v3

@@ 4 Datasets
To evaluate the defense potential of machine unlearning against MIA, we conduct experiments on four widely used datasets: CIFAR-10, Machine Unlearning Faces (MuFac), Purchase, and Texas100. The first two datasets are from the image domain, with Purchase and Texas100 be in tabular format. The four datasets represent sensitive real-world information, allowing us to assess the generalization and effectiveness of unlearning techniques in mitigating MIA across different domains. Below, we provide a detailed description of each dataset and its relevance to our experiments on machine unlearning.
@@ 4.1 Image Data
The CIFAR-10 dataset cifar10 is a popular benchmark dataset used in ML for image classification tasks. It consists of 60,000 color images, each of 32x32 pixels, categorized into 10 different classes. These classes are: airplanes, cars, birds, cats, deer, dogs, frogs, horses, ships, and trucks. Each class has 6,000 images, with 50,000 images designated for training and 10,000 for testing. The dataset is widely used due to its diversity and relatively small image size, making it a suitable challenge for various image recognition models, particularly in deep learning.
The MuFac Dataset choi2023towards is a multi-class age classification dataset, featuring over 13,000 Asian facial images with annotations for age groups and personal identities, ideal for machine unlearning research. The dataset provides a comprehensive collection of facial images with multiple factors influencing the appearance, including different angles, accessories (like glasses or hats), and emotions.
@@ 4.2 Tabular Data
The Purchase-100 dataset purchase100 is a customer purchase dataset widely used in privacy and adversarial ML research. It consists of 100 product classes, with each class representing a different product category. The dataset contains 600 binary features for each customer, indicating whether the customer purchased a specific product within a category. It is often used to study MIAs, making it useful for evaluating the privacy of ML models.
The Texas-100 dataset shokri2017membership , is a hospital discharge dataset that contains data from Texas hospitals. It includes 100 classes representing different diagnosis categories, with over 6,000 features per record. The features include patient demographics, treatment details and other hospital-related data. The Texas-100 dataset is commonly used in privacy research because of the sensitive nature of the information it contains, such as patient discharge status and procedures performed.
@@ 5 Experiments
In this section, we present the experimental setup and results of our study on the defense potential of machine unlearning algorithms against MIAs. Our experiments are designed to assess the unlearning performance of various state-of-the-art algorithms and evaluate their impact on different ML model architectures. We explore how effectively unlearning algorithms remove the influence of specific data points from trained models and how this influences the vulnerability of the models to MIAs. We also evaluate the utility of the models post-unlearning, ensuring that the removal of data does not significantly degrade predictive performance.
5.1 ML Model
All models were trained using an 80-20 train-test split. For each dataset and to evaluate the trained models against MIA, the target and shadow models used identical configurations, using five shadow models. Both target and shadow models used Multilayer Perceptron (MLP) architectures, for all datasets.
The training data for both the target and shadow models was divided equally, with a 50-50 split between target and shadow training data. MIAs were conducted using attack models that had equal proportions of members and non-members. The unlearning experiments followed the same data splitting as the training process, with an additional split to differentiate the retain and forget set (80-20 retain-forget split).
Table 1 summarizes the architectures used for each dataset, with MLP models employed across all datasets.
Each MLP architecture was specifically designed to balance efficiency and performance. The two fully connected (FC) layers with 512 and 256 units ensured that the models could handle the complexity of the image and tabular data while maintaining computational efficiency. The unlearning process, whether on specific classes or randomly selected rows, was designed to minimize the impact on retained data, ensuring that the models continued to perform well after unlearning.
@@ 5.1 ML Model
Dataset Epochs Learning Rate Layers and Units CIFAR-10 30 0.001 3 FC layers (512, 256, 128 units) MuFac 60 0.0007 3 FC layers (512, 256, 128 units) Purchase-100 100 0.01 1 FC layers (128 units) Texas-100 100 0.01 1 FC layers (128 units)
Table 1: ML Model Architectures
@@ 5.2 Unlearning Algorithms
In this paper, we used several machine unlearning methods, and we compared their capacity to address MIAs. Our goal is to achieve a better understanding of these algorithms, while trying to understand the potential gaps and strengths of unlearning algorithms against this type of attacks.
We include the SFTC algorithm perifanis2024sftc as our main focus, given its novel approach to balancing utility preservation and targeted forgetting. In addition we use NegGrad, the most straightforward baseline which modifies model parameters by applying gradient ascent on the forget set, aiming to erase the influence of the forget set. We also consider SCRUB, a teacher-student approach, which minimizes the divergence between the student and original models on the retain set while maximizing divergence on the forget set. Each algorithm is tested under identical conditions to ensure a fair comparison of their unlearning and defense capabilities.

### v4

@@ 5 Datasets
To evaluate the defense potential of machine unlearning against Membership Inference Attacks (MIAs), we conduct experiments on four widely utilized datasets: CIFAR-10, Machine Unlearning Faces (MuFac), Purchase-100, and Texas-100. These datasets offer a diverse representational scope, with CIFAR-10 and MuFac representing the image domain, while Purchase-100 and Texas-100 consist of tabular data.
The Purchase and Texas datasets are benchmarks within the privacy domain; they are integrated into frameworks such as MLPrivacyMeter Murakonda and Shokri (2020) and were utilized in the seminal work that introduced MIAs Shokri et al. (2017). Furthermore, these datasets, along with versions of CIFAR, are frequently employed in broader membership privacy research Nasr et al. (2018). Similarly, the MuFac dataset has emerged as a standard benchmark specifically for machine unlearning experiments Perifanis et al. (2024) Spartalis et al. (2025). Below, we provide a description of each dataset and its specific relevance to our evaluation of unlearning efficacy.
@@ 5.1 Image Data
The CIFAR-10 dataset Krizhevsky and Hinton (2009) is a foundational benchmark for image classification, consisting of 60,000 32x32 color images across 10 distinct classes. While the dataset’s categories are general, it serves as a critical proxy for user-contributed image repositories where individuals may assert their ’right to be forgotten.’ In this context, each image represents a unique data contribution that a model may inadvertently memorize during training.
Evaluating unlearning on CIFAR-10 allows us to measure how effectively an algorithm can remove the specific visual signature of a data point—preventing an adversary from confirming its presence in the training set. Furthermore, as a widely adopted baseline in privacy research, its use ensures that our evaluation of an algorithm’s ability to mitigate fundamental memorization is both reproducible and directly comparable with other methods.
The MuFac Dataset Choi and Na (2023) is a large-scale collection featuring over 13,000 Asian facial images. The dataset is comprehensive, with images capturing multiple factors influencing appearance, such as different angles, accessories (like glasses or hats), and emotional expressions. While the data is annotated with both age groups and personal identities, we clarify that the specific classification task used in our experiments is multi-class age classification.
The MuFac dataset serves as a high-fidelity proxy for privacy-sensitive biometric systems, where the right to be forgotten’ is particularly critical due to the immutable nature of facial features. Its detailed annotations, specifically the explicit link between individual images and personal identities, provide a rigorous framework for evaluating an algorithm’s ability to erase a specific biometric fingerprint. By targeting individual identities for removal while maintaining the model’s capacity to classify age for the remaining population, we can simulate real-world requirements for facial analysis technologies. This ensures that a model can comply with a user’s deletion request by successfully masking their membership traces without compromising the system’s overall functional utility.
@@ 5.2 Tabular Data
The Purchase-100 dataset Shokri et al. (2011), derived from Kaggle’s “Acquire Valued Shoppers Challenge”, consists of 197,324 customer records represented by 600-dimensional binary feature vectors. Each vector indicates whether a customer purchased products from specific category-company pairs. The classification task is to predict which of 100 behavioral segments—derived through clustering similar purchasing patterns, a given customer belongs to.
Purchase-100 serves as a critical benchmark for machine unlearning due to the high sensitivity of the transaction categories, which include pharmaceuticals, health products, and lifestyle choices. These features can reveal intimate details about an individual’s health conditions and personal circumstances. Despite the dataset’s overall size, the high-dimensional and sparse nature of the features makes individual purchase patterns potentially unique. This creates a realistic scenario where organizations must comply with deletion requests to mask these sensitive behavioral signatures, balancing the need for accurate market segmentation with the obligation to prevent adversarial re-identification through MIAs.
The Texas-100 dataset Shokri et al. (2017) is a hospital discharge dataset consisting of patient records from Texas hospitals. Each record contains over 6,000 features, including demographics, treatment details, and hospital-related metadata. The classification task in our experiments is to predict the patient’s primary diagnosis across 100 distinct categories.
Due to the presence of highly sensitive information, such as discharge status and surgical procedures, Texas-100 is a standard benchmark for privacy-preserving machine learning. Its high dimensionality makes individual patient records highly unique, creating a significant risk for membership inference. In the context of machine unlearning, this dataset represents a critical healthcare application where a model must be able to erase specific clinical fingerprints upon request to prevent the unauthorized disclosure of a patient’s medical history, all while maintaining high accuracy for general diagnostic research.
@@ 6 Experiments
In this section, we present the experimental setup and results of our study regarding the defense potential of machine unlearning algorithms against Membership Inference Attacks (MIAs). Our experiments are designed to systematically evaluate the performance of state-of-the-art unlearning algorithms across diverse model architectures. Specifically, we investigate how effectively these algorithms remove the influence of target data points while monitoring the resulting changes in the model’s vulnerability to MIAs.
The experiments were conducted on a workstation running Ubuntu 24.04.3 LTS, equipped with an AMD Ryzen™ 7 6800HS processor (16 logical cores) and an AMD Radeon™ 680M integrated graphics unit. For accelerated model training and unlearning, we utilized an NVIDIA GeForce RTX™ 3050 Laptop GPU (4GB VRAM) and a 512.1 GB SSD for high-speed data access and storage. All unlearning procedures were implemented using Python-based deep learning frameworks, ensuring that hardware utilization was optimized for the high-dimensional datasets described previously.
6.1 Target Model Methodology
The total training dataset was bifurcated equally using a 50-50 split between the target and shadow training sets. This partitioning strategy ensures that the shadow models are trained on data drawn from the same underlying distribution as the target model, while remaining strictly disjoint from the target model’s actual training samples.
All target models were trained using an 80/20/20 train-validation-test split, derived from the 50% of the total dataset allocated to the target model’s environment. We implemented various architectures for each dataset to observe how different levels of model complexity influence both predictive capacity and the ’shielding’ efficacy provided by unlearning against Membership Inference Attacks (MIAs).
In this study, we employed a spectrum of target models, ranging from simple baselines to complex, real-world architectures. Table 1 summarizes the architectures used for each dataset, with Multi-Layer Perceptrons (MLPs) serving as a common baseline across all data types. We categorized these models into three groups: simple, intermediate, and complex architectures. In this classification, the TargetModelID denotes the complexity group, while the suffixes a, b, c, and d correspond to the CIFAR-10, MUFAC, Purchase-100, and Texas-100 datasets, respectively.
The evaluated models include standard fully connected neural networks, Convolutional Neural Networks (CNNs), and more complex residual architectures such as ResNet. ResNet He et al. (2016) a deep CNN architecture that utilizes “skip connections” (or shortcut connections) within its residual blocks. These connections allow gradients to bypass certain layers, effectively mitigating the vanishing gradient problem and enabling the training of extremely deep networks. By learning residual mappings rather than direct transformations, ResNet facilitates the optimization of deep models, achieving superior performance in computer vision tasks such as image classification.
In all experiments, the target models were trained using hyperparameters that yielded the optimal test accuracy. To ensure the validity and fairness of our comparative analysis, we utilized the Optuna Python library Akiba et al. (2019), which employs a Tree-structured Parzen Estimator (TPE) to tune key parameters, including the number of hidden-layer neurons, the number of layers, the learning rate, the epoch count, and the batch size. Once the optimal hyperparameters were identified, the models were deliberately trained for additional epochs to induce an approximate 30% generalization gap between the training and test accuracy. This controlled overfitting was implemented to increase model susceptibility to attacks, thereby allowing us to evaluate whether unlearning algorithms can effectively protect even the most vulnerable, worst-case models.
6.2 Membership Inference Attack Configuration
6.2.1 Attack Model
For the implementation of the Attack Model of the MIA, we utilized a Random Forest classifier as the meta-model (attack model). The classifier was configured with 100 estimators to ensure a robust and stable decision boundary. The selection of this architecture was motivated by the following factors:
  • $\bullet$
    Non-linear Relationship Mapping: The prediction vectors (softmax probabilities) from target models often exhibit complex, non-linear distributions. Random Forests are inherently capable of capturing these high-dimensional interactions without requiring feature scaling or assuming a specific underlying distribution.
  • $\bullet$
    Variance Reduction via Bagging: By aggregating the predictions of 100 individual decision trees through bootstrap aggregating (bagging), the model significantly reduces variance. This prevents the attack model from overfitting to specific artifacts of the shadow models, ensuring better generalization when attacking the unseen target model.
  • $\bullet$
    Dimensionality Handling: In datasets such as Purchase-100 or Texas-100, where the output vector consists of 100 classes, the attack model must process high-dimensional input. Random Forests handle multi-modal probability scores efficiently, identifying subtle differences in confidence levels between members and non-members.
  • $\bullet$
    Stability and Efficiency: The choice of 100 estimators provides an optimal balance between computational overhead and predictive stability.
6.2.2 Shadow Model
For the shadow model architecture, as proposed by Shokri et al. (2017), we utilized the same architecture as the target model in each scenario. This approach assumes prior knowledge of the target model’s architecture by the adversary.While assuming knowledge of the target architecture represents a ’gray-box’ framing, this choice is justified by several real-world considerations. First, in the modern ML ecosystem, many organizations utilize standardized, state-of-the-art architectures (e.g., ResNet, VGG, or standard MLP bottlenecks) which are easily guessable by an informed adversary. Second, a system’s security should not rely on the secrecy of the algorithm but on the secrecy of the data or keys; therefore, evaluating the ’worst-case’ scenario where the attacker knows the architecture provides a more robust upper bound of the privacy risk. Finally, research has shown that Membership Inference Attacks are effective even when there is an architectural mismatch Zou et al. (2020), but using the same architecture ensures that the observed privacy leakage, ot lack of leackage, is attributable to the unlearning algorithm’s performance rather than structural differences between models. In our experiments we used 5 shadow models.
@@ 6.2.2 Shadow Model
TargetModelID Dataset Epochs Learning Rate Layers and Units 1a CIFAR-10 30 0.001 3 FC layers (512, 256, 128 units) 1b MuFac 60 0.0007 3 FC layers (512, 256, 128 units) 1c Purchase-100 100 0.01 1 FC layers (128 units) 1d Texas-100 100 0.01 1 FC layers (128 units) 2a CIFAR-10 30 0.001 2 Conv Blocks (16, 32 filters)+1FC layers (128 units) 2b MuFac 60 0.0007 2 Conv Blocks (16, 32 filters)+1FC layers (128 units) 2c Purchase-100 100 0.01 2 FC layers (256, 128 units) 2d Texas-100 100 0.01 2 FC layers (256, 128 units) 3a CIFAR-10 30 0.001 Resnet 3b MuFac 60 0.0007 Resnet 3c Purchase-100 100 0.01 4-Layer MLP (1024, 512, 256 units) + Dropout 3d Texas-100 100 0.01 4-Layer MLP (512, 128, 128 units) + Dropout
Table 1: Target Model Architectures
6.3 Overfitting as a Worst-Case Privacy Benchmark
A notable characteristic of our experimental setup is the significant generalization gap observed in the target models, which reaches up to 30%. While such levels of overfitting are unrepresentative of high-performing production models, this setting was intentionally chosen to maximize the Membership Inference (MI) signal.
Since MIA success is fundamentally bounded by the degree of empirical risk minimization and memorization, utilizing overfitted models serves as a “stress test” for the unlearning algorithms under evaluation. By amplifying the signal that the unlearning process is intended to erase, we can more precisely measure the privacy residual, the traces of training data that remain even after an unlearning request is processed.
This ’worst-case’ framing establishes a conservative upper bound on privacy risk ; if an algorithm can successfully mitigate a strong membership signal in an overfitted model, its defensive efficacy is expected to be even more robust in well-regularized production environments where the baseline signal is naturally weaker. Furthermore, this controlled gap ensures that the observed privacy results are attributable to the unlearning algorithm’s performance rather than a failure of the attack to detect members due to low baseline memorization.
@@ 6.4 Unlearning Algorithms
In this study, we focus on a comparative evaluation of three representative machine unlearning algorithms: NegGrad Golatkar et al. (2020) , SCRUB Kurmanji et al. (2023), and SFTC Perifanis et al. (2024). These methods were carefully selected to cover distinct and widely recognized families of unlearning approaches. NegGrad exemplifies gradient-based parameter manipulation techniques aimed at directly removing the influence of target data. SCRUB uses a teacher-student framework that leverages knowledge distillation and divergence maximization to balance retention and unlearning. SFTC, a recent state-of-the-art method, integrates selective fine-tuning with targeted confusion to optimize unlearning effectiveness. Together, these algorithms represent a diverse spectrum of strategies addressing the unlearning challenge. While many machine unlearning algorithms exist in the literature, incorporating dozens of variants and hybrids, this study focuses on these three due to their strong empirical performance and methodological variety, which allows for a meaningful yet tractable comparison consistent with practices in recent benchmarking and survey works. This approach aligns with the field’s current standards, where in-depth analysis typically centers on a carefully chosen subset of leading algorithms to provide clear insights and rigorous evaluation.
We include the SFTC algorithm Perifanis et al. (2024) as our main focus, given its novel approach to balancing utility preservation and targeted forgetting. In addition we use NegGrad, the most straightforward baseline which modifies model parameters by applying gradient ascent on the forget set, aiming to erase the influence of the forget set. We also consider SCRUB, a teacher-student approach, which minimizes the divergence between the student and original models on the retain set while maximizing divergence on the forget set. Each algorithm is tested under similar conditions to ensure a fair comparison of their unlearning and defense capabilities.
6.5 Experimenting Methodology
In these experiments, we apply unlearning algorithms to the target models and perform an MIA at each epoch to measure both MIA Forget Accuracy and MIA Retain Accuracy. This approach provides a high-granularity estimation of the attack’s efficacy throughout the unlearning process, enabling a deeper understanding of the temporal unlearning dynamics.
Our primary objective is to observe whether the MIA accuracy on the forget set diminishes over time keeping the MIA Retain accuracy stable. Ideally, a model should unlearn specific data such that it yields a low MIA forget accuracy, behaving as if those samples had never been part of the training set. Simultaneously, it is essential to maintain a stable MIA retain accuracy to ensure that the model’s knowledge of the remaining data is preserved. A decrease in MIA retain accuracy may indicate that the unlearning method is starting to understand after unlearning the Retain Data with a different way. Conversely, a significant increase in retain accuracy may suggest that the retain set is becoming more vulnerable to privacy leaks as the unlearning process progresses.
Maintaining a stable MIA Retain Accuracy alongside a significantly diminished MIA Forget Accuracy indicates that the unlearning algorithm has the potential to effectively mitigate privacy risks for the forgotten cohort without compromising the privacy profile of the remaining data.
To ensure a rigorous and comparative analysis, a uniform data partitioning strategy was applied across all evaluated algorithms. Specifically, 20% of the training samples were allocated to the forget set, while the remaining 80% served as the retain set.